OpenStudy (anonymous):

Find the coordinates of the midpoint of the segment whose end points are H(8,2) and K (6,10) ? (2,8) (1,4) (7,6) (14,12) Anyone know the formula so I can try and work it out first ?

OpenStudy (anonymous):

take the average in each coordinate aka add and divide by 2

OpenStudy (anonymous):

in math teacher speak \[\left(\frac{x_1+x_2}{2},\frac{y_1+y_2}{2}\right)\]

OpenStudy (anonymous):

I got (7,6) :)?

OpenStudy (anonymous):

oui as usual it is C